OpenStudy (anonymous):

A random number generator picks a number from 1 to 9 in a uniform manner. Find the 90th percentile. (Round your answer to one decimal place.) My set up is (k-9)(1/8)=0.90 why is this wrong

OpenStudy (nincompoop):

why did you use that "set up?"

OpenStudy (anonymous):

in my book it says to use it

OpenStudy (anonymous):

(base)(height)=0.90

OpenStudy (nincompoop):

first, what do you know about percentile rank in your own words

OpenStudy (anonymous):

isn't it like how many # were chosen...90%?

OpenStudy (nincompoop):

so you're being asked for the number that ranks 90th percentile

OpenStudy (anonymous):

right

OpenStudy (nincompoop):

how would you look for that number

OpenStudy (anonymous):

that's where I am getting confused because my set up is wrong? would it be like 90%..of 9

OpenStudy (nincompoop):

let us list your numbers

OpenStudy (nincompoop):

is this how you have it set up? 1, 2, 3, 4, 5, 6, 7, 8, 9

OpenStudy (anonymous):

okay yes

OpenStudy (nincompoop):

there are two ways to get percentile ranks one is: \(\sf \huge \frac{B\times (0.5E)}{n} \times 100 \) B: scores below x E: number of scores equal x n: total number the other is: \(\sf \huge \frac{B}{n} \times 100 \)

OpenStudy (nincompoop):

if you're going to use the second one, the set up should be \(\sf \large \frac{B}{9} \times 100 = 90 \) remember that your B will be lower than the actual score you're looking for, but let us focus on properly solving for B first

OpenStudy (anonymous):

so it would be 8.1?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

I thought it ment to keep the .1

OpenStudy (nincompoop):

round off is the same as estimation and you need to learn about place values.

OpenStudy (nincompoop):

9.1 rounded off to 1 decimal means you focus on the first decimal if it is between 0-4 then it becomes 0 if it is between 5-9 then you add 1 to the number that preceded it 9.1 = 9 9.5 = 10 9.67 = 10 etcetera etcetera

OpenStudy (nincompoop):

let us check if our answer makes sense http://www.wolframalpha.com/input/?i=8%2F9

OpenStudy (anonymous):

I understand that, but the question says "Round your answer to one decimal place" so therefore the answer would be like "9.4" keeping the 4

OpenStudy (nincompoop):

8 out 9 makes almost 90%

OpenStudy (nincompoop):

no

OpenStudy (anonymous):

when I practice another problem with numbers 2-9 and asking for the 80th percentile, the answer is 7.6 ??

OpenStudy (anonymous):

the .6 is kept

OpenStudy (nincompoop):

http://www.mathsisfun.com/rounding-numbers.html

OpenStudy (nincompoop):

you're right you keep it

OpenStudy (anonymous):

so youre saying when it is smaller than .4 it is rounded to 0

OpenStudy (nincompoop):

but 1, 2, 3, 4, 5, 6, 7, 8, 9 tell me where is 8.1

OpenStudy (anonymous):

so you're saying the answer is 8

OpenStudy (nincompoop):

in scores 1 - 9 and if they are just 1, 2 .... 9 then yes unless someone actually has a score of 8.1 then it is 8.1

OpenStudy (anonymous):

so I tried 8.1 and 8 and it is saying it is wrong

OpenStudy (nincompoop):

how about the first formula

OpenStudy (anonymous):

the one that I used? nope

OpenStudy (nincompoop):

I made a typo \(\huge \frac{(B + 0.5E )}{n} \times 100 \)

OpenStudy (nincompoop):

so 8.1 and 8.0 are wrong huh?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

did you put 8.0 with point?

OpenStudy (anonymous):

no I just did 8

OpenStudy (nincompoop):

try 8.0 because of the 10th place value

OpenStudy (nincompoop):

did it work?

OpenStudy (anonymous):

wait but I only have one try left so I would like to be really sure!

OpenStudy (anonymous):

the book says 8.2??

OpenStudy (nincompoop):

I don't know what you just did in the end, kropot but I do want to know why 8.2 because I want to learn if there's any other methods of solving this.

OpenStudy (anonymous):

I am really not sure. They give me an example of "random number generator picks a number from 2 to 9 in a uniform manner. Find the 80th percentile(Round your answer to one decimal place . and the answer is 7.6 I am trying to work backwards from the answer but I still do not know

OpenStudy (dan815):

(7.6-2)/(9-2)=0.8

OpenStudy (dan815):

using same method, (k-1)/(9-1)=0.9 k=?

OpenStudy (dan815):

k=8.2

OpenStudy (anonymous):

ohohoh I see

OpenStudy (anonymous):

thank you so much for this explanation!!!!

OpenStudy (anonymous):

what was the equation you used tho?

OpenStudy (dan815):

i looked at the example they gave you

OpenStudy (dan815):

basically what they are doing is

OpenStudy (dan815):